# Interpreting the power spectral density of a fluctuating colloidal current

### v1

@@ Abstract
The transport of molecules through biological and synthetic nanopores is governed by multiple stochastic processes that lead to noisy, fluctuating currents. Disentangling the characteristics of different noise-generating mechanisms is thus central to better understanding molecular transport at a fundamental level. Here, we study current noise experimentally at the single particle level by imaging colloidal particles driven through microfluidic channels by a difference in fluid pressure. In this scenario, currents fluctuate due to the random arrival times of particles into the channel and the distribution of particle speeds within the channel. We find that this results in a characteristic form of the power spectral density of the current, scaling as $\sim f^{0}$, at low frequencies and $\sim 1/f^{2}$ at high frequencies. To rationalise these scalings, we extend a model for shot noise with a finite transit time, borrowed from electronic circuit theory, to include the experimental distribution of transit times and compare this model to the experimental spectra. We find excellent agreement between the data and model across a range of driving pressures within an advection-dominated regime, thereby establishing concrete links between the power spectra scalings and underlying mechanisms for this experimental system. This paves the way for establishing a more systematic understanding of the links between features of power spectra and underlying molecular mechanisms in driven systems.
@@ I Introduction
The ability to sensitively control and manipulate the passage of macromolecules, molecules and ions through nanoscopic pores in a membrane is central to the function of many important emerging technologies, from biomedical sensors Howorka and Siwy (2009) to desalination devices Werber, Osuji, and Elimelech (2016). Here, biological membrane proteins represent a gold standard, combining high throughput with exquisite selectivity for specific molecular species Park et al. (2017). While it is possible to build effective devices around (modified) biological nanopores Schmidt (2016), recent advances in nanoscale fabrication techniques have motivated efforts to replicate this functionality in synthetic nanopore systems Dekker (2007). Such solid state pores could be built to order and would offer key advantages, such as greater robustness and the potential to exploit exotic properties of non-biological materials like graphene Sahu and Zwolak (2019); Radha et al. (2016); Emmerich et al. (2022); Caglar et al. (2020). Despite significant efforts, the level of control over transport through synthetic pores remains some way off that of their biological analogues Robin and Bocquet (2023), in part due to many open questions surrounding the governing principles of confined transport processes.
@@
One characteristic aspect of transport through nanoscale pores is significant fluctuations in measured currents. Over certain frequency regimes, these fluctuations arise from dynamic processes at the molecular level, such as the thermal fluctuations of ions Marbach (2021); Knowles, Keyser, and Thorneywork (2020); Smeets et al. (2008); Bezrukov et al. (2000) or adsorption to the pore surface Knowles et al. (2021); Gravelle, Netz, and Bocquet (2019); Robin et al. (2023a). As such, interpreting fluctuations in these systems can provide significant information on transport at the molecular level beyond that available from the magnitude of the current alone. In experiments this is generally approached via interpretation of the power spectral density (PSD) of the current, with ongoing efforts to understand the nature of fluctuations in both biological and synthetic porous systems Fragasso, Schmid, and Dekker (2020). In spite of this, unambiguously linking measured fluctuations to underlying mechanisms is very challenging even for relatively simple scenarios. This is a consequence of the inherent complexity of molecular systems, particularly when under confinement, which introduces behaviours not seen in bulk. For instance, steric and hydrodynamic interparticle interactions change under confinement, becoming longer ranged and even non-decaying with distance Cui et al. (2004); Misiunas et al. (2015); Wei, Bechinger, and Leiderer (2000); Robin et al. (2023b). Pore geometry can lead to current rectification Powell et al. (2011); Aarts et al. (2022), with further changes in transport if the confining pore is not rigid but fluctuating Marbach, Dean, and Bocquet (2018). Moreover, interactions of ions and molecules with the confining surface lead to complex electroosmotic flows Firnkes et al. (2010); Mc Hugh, Andresen, and Keyser (2019); Gubbiotti et al. (2022), and adsorption itself has been shown to affect the spectral content of nanopore currents Gravelle, Netz, and Bocquet (2019); Knowles et al. (2021).
@@
Figure 1: Quantifying fluctuations in colloidal currents. (a) Schematic of the experimental set-up in which macroscopic reservoirs at different heights produce a pressure difference across the channel. (b) A typical image of the colloidal system (c) The current at two magnifications for a box with $L=64~{}\mu$m, showing the arrival/departure of individual particles on short time scales, and the form of the current over longer time scales. (d) The power spectral density calculated from the current in (c).
@@
From an experimental perspective, challenges involved in fabricating nanoscale pores with precisely controlled features also makes interpreting and comparing data from molecular level experiments difficult. In contrast, colloidal models allow for sensitive control of key physicochemical properties, including particle size and shape Sacanna et al. (2013), interparticle interactions Yethiraj and Blaaderen (2003); Royall, Leunissen, and van Blaaderen (2003), confining geometry Yang et al. (2017) and driving force Misiunas and Keyser (2019); Liot et al. (2018) in systems where single particle Franosch et al. (2011); Krapf et al. (2019) and many-body Mackay et al. (2023) fluctuations can be visualised in detail. This has motivated various experimental studies on microscale transport processes in confinement, examining diffusive behaviour Siems et al. (2012); Yang et al. (2017); Dettmer et al. (2014); Thorneywork et al. (2020), the rheological properties of colloidal gels and glasses Isa, Besseling, and Poon (2007); Conrad and Lewis (2008), or driven flows exhibiting clogging and jamming Wyss et al. (2006); Marin et al. (2018); Souzy, Zuriguel, and Marin (2020). Fluctuations in colloidal transport are, however, comparatively less well explored.
@@
Here, we study fluctuations in the transport of colloidal particles driven through a microfluidic channel to determine how specific physical mechanisms in this system are manifested in the power spectral density. We first define an approach to calculate currents from individual particle trajectories and subsequently quantify fluctuations in the current via the power spectral density. Our power spectra exhibit a characteristic form, scaling as $\sim 1/f^{2}$ at high frequencies and exhibiting white noise, $\sim f^{0}$, at low frequencies. By considering relevant time scales for the particle transport, we link these scaling regimes to key noise-generating mechanisms in this system, namely the random arrival times of particles to the channel and the distribution of particle speeds within it. Finally, by drawing an analogy with noise in electronic circuits, we identify a suitable model for the fluctuations in our system as shot noise with a finite transit time and modify this to include the distribution of particle speeds within the channel. We compare spectra predicted from this model with our measured experimental spectra and find excellent agreement between the two.
@@ II Experimental methods
The experimental system has previously been reported Knowles et al. (2022) and is illustrated schematically in Fig. 1(a). This consists of a microfluidic chip designed to include two reservoirs linked by channels with length $L\sim 100~{}\mu$m, height $h=8~{}\mu$m and a width $14~{}\mu$m. The device is fabricated by replica moulding the main structure in poly-dimethylsiloxane (PDMS) and then plasma bonding this component onto a glass slide coated with a thin layer of PDMS. The chip is filled with a suspension of $\sigma=2.8~{}\mu$m carboxylate functionalised melamine formalydehyde particles. The high density of the colloidal particles with respect to the solvent causes them to sediment rapidly onto the base of the chip, forming a quasi-2D monolayer confined by gravity. Following assembly, the microfluidic device is attached to two macroscopic fluid reservoirs that impose a pressure difference across the chip to drive particles through the channels.
@@ III.1 Quantifying particle transport as a current
The pressure-driven fluid flow creates an advection dominated transport of particles inside the channels with particles primarily following streamlines parallel to the channel wallsKnowles et al. (2022). To characterise this transport as a current,
we define a region of interest (ROI) of length $L$ which spans the width of the channel, as illustrated in Fig. 1(b). When crossing the ROI each particle is defined to contribute a rectangular pulse to the current as:
@@
To reduce the statistical error on values of $S(f)$, we take the average of PSDs calculated from multiple short sections of the total current trace, using Welch’s method Press et al. (1992). The resulting spectra is smoothed further by logarithmically binning along the frequency axis and taking the average noise power in each bin.
@@
For driven transport, the most obvious time scale for the system is the mean time it takes a particle to cross the ROI. For this dataset, this time is $\sim 2$s, which is not dissimilar to the crossover frequency. As such, the high frequency regime mainly corresponds to times shorter than the mean time it takes a particle to cross the ROI, implying that the $1/f^{2}$ scaling and oscillations link to the particle transit through the channel. In contrast, the low frequency regime corresponds to behaviour over times longer than the particle transit times, and so instead reflects fluctuations in the current associated with the arrival times of the particles into the channel.
@@ III.2 Variation of spectra with ROI length $L$
To further investigate the crossover between the two regimes in the spectra, we next consider the variation in the PSD with the length of the ROI, $L$. Fig. 2(a) illustrates the qualitative differences between current traces measured in longer or shorter regions. A larger value of $L$ corresponds to a more smoothly varying current, comprised of many broad, overlapping pulses, whereas at smaller values of $L$ the discretisation of current carriers becomes more significant, leading to currents comprised of distinct spikes. While the apparent magnitude of the current increases as $L$ decreases, this is a consequence of dividing the pulse per particle by the transit time. This ensures that the current as a measure of the mean number of transported particles is independent of the ROI size, despite the fact that the number of particles instantaneously occupying a shorter ROI is lower due to the shorter transit times.
@@
The corresponding spectra are presented in Fig. 2(b). All spectra have the same form as that in Fig. 1(d) with the most notable difference between spectra being the shift in the corner frequency to lower values with increasing $L$. This supports the link between corner frequency and mean particle transit time as $f_{c}\sim 1/\tau$, as increasing the length of the ROI will increase the mean transit time.
For low frequencies, the spectra are essentially identical, with a plateau value of approximately double the mean current through the ROI. This is consistent with the picture of a low frequency regime reflecting fluctuations in the arrival of particles into the ROI; the mean current depends on the total number of particles that pass through any region of the channel which is equal to the mean arrival rate due to particle number conservation. The coincidence of the fine structure of the PSD at low frequencies is a consequence of the fact that the current in a larger ROI can be expressed in terms of the current through any smaller ROI within the same channel. This is discussed in Appendix A.
@@ III.3 Modelling shot noise with a finite transit time
Having established the phenomenological behaviour of the power spectral density of our system in different frequency regimes we now seek to more quantitatively link features of the PSD to molecular mechanisms. At low frequencies. the dominant mechanism appears to be the small number noise, or ‘shot noise’, associated with random arrival of particles into the channel. This type of noise has been widely studied and extensively characterised in the context of electronic circuits MacDonald (2006); Kogan (1996). As such, these works provide a starting point for modelling the current in our colloidal system.
@@
for an average current, $I_{0}$, and elementary charge, $e$. The root mean squared noise, $\sigma$, measured in some finite bandwidth, $\Delta f$, is then
$\sigma=(2eI_{0}\Delta f)^{1/2}.$ (5)
As the signal to noise ratio for purely shot noise varies with $I_{0}^{-1/2}$, shot noise can be neglected for large currents but is significant for situations where there are small numbers of particles making up the current as in this study.
Pure shot noise assumes that pulses contributing to the current are instantaneous. Yet in many physical systems, pulses have a finite duration, for example due to the time required for a charge carrier to move from emitter to detector. This is the scenario for the colloidal model considered here, as pulses have a finite duration equal to the time for particles to cross the ROI (see Eq. (III.1)). The effect of finite transit times on shot noise can be modelled analytically, and in the following we outline the derivation from MacDonald MacDonald (2006).
Figure 3: The PSD of shot noise with a single, finite transit time $\tau$ as described by Eq. 7. The maximum magnitude of the spectra has been normalised to 1, and the frequency has been rescaled by the characteristic frequency $1/\tau$.
If we assume that all particles cross the channel with the same velocity, every particle will contribute a rectangular pulse of length $\tau$ to the signal. The real-space autocorrelation function for the current fluctuations can then be expressed in terms of a single pulse, $i_{1}(t)$, starting at $t=0$ as:
$\psi(T)$ $=\langle I(t)I(t+T)\rangle$
[…]
$=\frac{I_{0}}{\tau}\left(1-\frac{T}{\tau}\right).$ (6)
[…]
leading to the final expression:
$S(f)=\frac{2I_{0}\sin^{2}(\pi f\tau)}{\pi^{2}f^{2}\tau^{2}}$ (7)
A plot of Eq. 7 is shown in Fig. 3, where the frequency has been rescaled by the characteristic frequency, $1/\tau$, and the magnitude is rescaled by $1/2I_{0}$. In line with the experimental results the spectra show the same two scaling regimes, separated by a corner frequency, $f_{c}=1/(2\tau)$. Notably, the PSD also shows characteristic oscillations at high frequencies with zeros at $f=\frac{n}{\tau}$ for integer values of $n$, although these oscillations are much more pronounced than in the experimental spectra in Fig. 1(d).

### v2

# Supplementarymaterial

@@ Abstract
The transport of molecules through biological and synthetic nanopores is governed by multiple stochastic processes that lead to noisy, fluctuating currents. Disentangling the characteristics of different noise-generating mechanisms is central to better understanding molecular transport at a fundamental level but is extremely challenging in molecular systems, due to their complexity and relative experimental inaccessibility. Here, we construct a colloidal model microfluidic system for the experimental measurement of particle currents, where the governing physical properties are directly controllable and particle dynamics directly observable, unlike in the molecular case. Currents of hard spheres fluctuate due to the random arrival times of particles into the channel and the distribution of particle speeds within the channel, which results in characteristic scalings in the power spectral density. We rationalise these scalings by quantitatively comparing to a model for shot noise with a finite transit time, extended to include the distribution of particle speeds. Particle velocity distributions sensitively reflect the confining geometry and we interpret and model these in terms of the underlying fluid flow profiles. Finally, we explore the extent to which details of these distributions govern the form of the resulting power spectral density, thereby establishing concrete links between the power spectral density and underlying mechanisms for this experimental system. This paves the way for establishing a more systematic understanding of the links between characteristics of transport fluctuations and underlying molecular mechanisms in driven systems such as nanopores.
@@ I Introduction
The ability to sensitively control and manipulate the passage of macromolecules, molecules and ions through nanoscopic pores in a membrane is central to the function of many important emerging technologies, from biomedical sensors Howorka2009 to desalination devices Werber2016 . Here, biological membrane proteins represent a gold standard, combining high throughput with exquisite selectivity for specific molecular species Park2017 . While it is possible to build effective devices around (modified) biological nanopores Schmidt2016 , recent advances in nanoscale fabrication techniques have motivated efforts to replicate this functionality in synthetic nanopore systems Dekker2007 . Such solid state synthetic pores could be built to order and would offer key advantages, such as greater robustness and the potential to exploit exotic properties of non-biological materials like graphene Sahu2019 ; Radha2016 ; Emmerich2022 ; Caglar2020 . Despite significant efforts, the level of control over transport through synthetic pores remains some way off that of their biological analogues Robin2023review , in part due to many open questions surrounding the governing principles of confined transport processes.
@@
One characteristic aspect of transport through nanoscale pores is significant fluctuations in measured currents. Over certain frequency regimes, these fluctuations arise from dynamic processes at the molecular level, such as the thermal fluctuations of ions Marbach2021 ; Knowles2020 ; Smeets2008 ; Bezrukov2000 or adsorption to the pore surface Knowles2021 ; Gravelle2019 ; Robin2023 . As such, interpreting fluctuations in these systems can provide significant information on transport at the molecular level beyond that available from the magnitude of the current alone. In experiments this is generally approached via interpretation of the power spectral density (PSD) of the current, with ongoing efforts to understand features of the PSD in both biological and synthetic porous systems Fragasso2020 . In spite of this, unambiguously linking measured fluctuations to underlying mechanisms is very challenging even for relatively simple scenarios. This is a consequence of the inherent complexity of molecular systems, particularly when under confinement, which introduces behaviours not seen in bulk. For instance, steric and hydrodynamic interparticle interactions change under confinement, becoming longer ranged and even non-decaying with distance Cui2004 ; Misiunas2015 ; Wei2000 ; Robin2023b . Pore geometry can lead to current rectification Powell2011 ; Aarts2022 , with further changes in transport if the confining pore is not rigid but fluctuating Marbach2018 . Interactions of ions and molecules with the confining surface lead to complex electroosmotic flows Firnkes2010 ; McHugh2019 ; Gubbiotti2022 , and adsorption itself has been shown to affect the spectral content of nanopore currents Gravelle2019 ; Knowles2021 . Moreover, as the confinement is heterogeneous, with finite length pores connecting wider reservoirs, long time fluctuations of the system may primarily depend on how behaviour in the reservoir changes pore entrance statistics Robin2023 ; Fragasso2019 .
@@
Figure 1: Quantifying fluctuations in colloidal currents. (a) Schematic of the experimental set-up in which macroscopic reservoirs at different heights produce a pressure difference across the channel. (b) A typical image of the colloidal system. (c) The current at two magnifications for a box with $L=64~{}\mu$m, showing the arrival/departure of individual particles on short time scales, and the form of the current over longer time scales. (d) The power spectral density calculated from the current in (c).
@@
From an experimental perspective, challenges involved in fabricating nanoscale pores with precisely controlled features also makes interpreting and comparing data from molecular level experiments difficult. In contrast, colloidal models allow for sensitive control of key physicochemical properties, including particle size and shape, interparticle interactions, confining geometry Yang2017 and driving force Misiunas2019 ; Liot2018 in systems where single particle Franosch2011 ; Krapf2019 and many-body Mackay2023 fluctuations can be visualised in detail. While there are obvious physical differences between molecular and colloidal systems, analogous physical processes take place at both length scales.Thorneywork2020 This has motivated various experimental studies on microscale transport processes in confinement, examining diffusive behaviour Siems2012 ; Yang2017 ; Dettmer2014 ; Thorneywork2020 , or driven flows exhibiting clogging and jamming Wyss2006 ; Marin2018 ; Souzy2020 . Yet, colloidal systems are typically not set up to address directly analogous questions to those in existing nanopore experiments and specifically have rarely considered current fluctuations in the frequency domain.
@@
Here, we show that fluctuations in currents of colloidal hard spheres in an experimental model ‘micropore’ can be fully characterised and unambiguously interpreted with respect to underlying physical properties of the system. By measuring the current from individual particle trajectories, we take advantage of direct experimental information inaccessible at the nanoscale. This allows us to quantitatively understand fluctuation mechanisms that are otherwise difficult to probe directly in molecular experiments. We first define an approach to calculate currents from individual particle trajectories and subsequently quantify fluctuations in the current via the power spectral density. Our power spectra exhibit a characteristic form, scaling as $\sim 1/f^{2}$ at high frequencies and exhibiting white noise, $\sim f^{0}$, at low frequencies. By considering relevant time scales for the particle transport, we link these scaling regimes to key noise-generating mechanisms in this system, namely the random arrival times of particles to the channel and the distribution of particle speeds within it. We then demonstrate quantitative agreement between our data and a model for shot noise with a finite transit time, extended to include the distribution of particle speeds. Particle velocity distributions sensitively reflect the confining geometry and we interpret and model these in terms of the underlying fluid flow profiles. Finally, we explore the extent to which details of these distributions govern the form of the resulting power spectral density, thereby establishing concrete links between the power spectral density and underlying mechanisms for this experimental system.
@@ II Experimental methods
The experimental system has previously been reported Knowles2022 and is illustrated schematically in Fig. 1(a). This consists of a microfluidic chip designed to include two reservoirs linked by channels with length $L\sim 100~{}\mu$m, height $h=8~{}\mu$m and a width $\sim 14~{}\mu$m. The device is fabricated by replica moulding the channel structure in poly-dimethylsiloxane (PDMS) and then plasma bonding this component onto a glass slide coated with a thin layer of PDMS. The chip is filled with a suspension of $\sigma=2.8~{}\mu$m carboxylate functionalised melamine formalydehyde particles. The high density of the colloidal particles with respect to the solvent causes them to sediment rapidly onto the base of the chip, forming a quasi-2D monolayer confined by gravity. Following assembly, the microfluidic device is attached to two macroscopic fluid reservoirs that impose a pressure difference across the chip to drive particles through the channels.
@@ III.1 Quantifying particle transport as a current
The pressure-driven fluid flow creates an advection dominated transport of particles inside the channels with particles primarily following streamlines parallel to the channel wallsKnowles2022 . To characterise this transport as a current, we define a region of interest (ROI) of length $L$ which spans the width of the channel, as illustrated in Fig. 1(b). When crossing the ROI each particle is defined to contribute a rectangular pulse to the current as:
@@
To reduce the statistical error on values of $S(f)$, we take the average of PSDs calculated from multiple short sections of the total current trace, using Welch’s method Press1992 . The resulting spectra are smoothed further by logarithmically binning along the frequency axis and taking the average noise power in each bin.
@@
For driven transport, the most obvious time scale for the system is the mean time it takes a particle to cross the ROI. For this dataset, this time is $\sim 2$ s, which is not dissimilar to the crossover frequency. As such, the high frequency regime mainly corresponds to times shorter than the mean time it takes a particle to cross the ROI, implying that the $1/f^{2}$ scaling and oscillations link to the particle transit through the channel. In contrast, the low frequency regime mainly corresponds to times longer than the mean time it takes a particle to cross the ROI. It instead reflects fluctuations associated with the transport of particles from the reservoir into the channel.
@@ III.2 Variation of spectra with ROI length $L$
To further investigate the crossover between the two regimes in the spectra, we next consider the variation in the PSD with the length of the ROI, $L$. Fig. 2(a) illustrates the qualitative differences between current traces measured in longer or shorter regions. A larger value of $L$ corresponds to a more smoothly varying current, comprised of many broad, overlapping pulses, whereas at smaller values of $L$ the instantaneous current is due to only one or two particles, leading to currents comprised of distinct spikes. These spikes are both narrower and taller than the pulses in longer ROIs, due to the shorter transit times. The mean current is, however, independent of the ROI size, as while fewer particles instantaneously occupy a shorter ROI, their transit times are also shorter.
@@
The corresponding spectra are presented in Fig. 2(b). All spectra have the same form as that in Fig. 1(d) with the most notable difference between spectra being the shift in the corner frequency to lower values with increasing $L$. This supports the link between corner frequency and mean particle transit time as $f_{c}\sim 1/\tau$, as increasing the length of the ROI will increase the mean transit time. For low frequencies, the spectra are essentially identical, with a plateau value of approximately double the mean current through the ROI. This is consistent with the picture of a low frequency regime reflecting fluctuations in the arrival of particles into the ROI; the mean current depends on the total number of particles that pass through any region of the channel which is equal to the mean arrival rate due to particle number conservation. The coincidence of the fine structure of the PSD at low frequencies is a consequence of the fact that the current in a larger ROI can be expressed in terms of the current through any smaller ROI within the same channel. This is discussed in Appendix A.
@@ III.3 Modelling shot noise with a finite transit time
Having established the phenomenological behaviour of the power spectral density of our system in different frequency regimes we now seek to more quantitatively link features of the PSD to physical mechanisms. At low frequencies, the dominant mechanism appears to be the small number noise, or ‘shot noise’, associated with random arrival of particles into the channel. This type of noise has been widely studied and extensively characterised in the context of electronic circuits MacDonald2006 ; Kogan1996 . As such, these works provide a starting point for modelling the current in our colloidal system.
@@
for an average current, $I_{0}$, and elementary charge, $e$. The root mean squared noise, $\Delta I_{rms}$, measured in some finite bandwidth, $\Delta f$, is then
$\Delta I_{rms}=(2eI_{0}\Delta f)^{1/2}.$
As the signal to noise ratio for purely shot noise varies with $I_{0}^{-1/2}$, shot noise can be neglected for large currents but is significant for situations where there are small numbers of particles making up the current, as in this study.
Figure 3: A comparison of the experimental PSD with models for shot noise with a finite transit time. (a) Experimental data (points) compared to a single, finite transit time model (line), as described by Eq. 5. (b)-(c) Experimental data (points) compared to the distributed transit time model of Eq. 6. Panels (a) and (b) show spectra with $L=64~{}\mu$m and mean velocity, $\langle v\rangle=36.1~{}\mu\rm{ms}^{-1}$, whilst panel (c) shows spectra with $L=64~{}\mu$m and $\langle v\rangle=6.8~{}\mu\rm{ms}^{-1}$. Insets show the corresponding experimental probability distribution of the velocity, $P(v)$.
Pure shot noise assumes that pulses contributing to the current are instantaneous. Yet in many physical systems, pulses have a finite duration, for example due to the time required for a charge carrier to move from emitter to detector. This is the scenario for the colloidal model considered here, as pulses have a finite duration equal to the time for particles to cross the ROI (see Eq. III.1). The effect of finite transit times on shot noise can be modelled analytically (see Appendix B) leading to an expression for $S(f)$ as
$S(f)=\frac{2I_{0}\sin^{2}(\pi f\tau)}{\pi^{2}f^{2}\tau^{2}}.$ (5)
In Fig. 3(a) we show a comparison between a typical experimental spectrum and the prediction of Eq. 5 using the experimental mean current as $I_{0}$ and using the mean particle velocity in the channel to determine the single transit time, $\tau$. The modelled spectra show the same two scaling regimes, separated by a corner frequency, $f_{c}=1/(2\tau)$ and both the measured and modelled spectra show characteristic oscillations at high frequencies. Oscillations in the model are much more pronounced than in the experimental spectra, however, and so agreement at high frequencies is lacking.